# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ Appendix A Filtering
Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection. The pipeline proceeds in three steps. A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales. A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds=\Sigma^{XX,i}(t)$. A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
@@ A.2 Specialization to the physical observation model
$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ in the general framework of Proposition A.2.
@@
Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
@@
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$. Thus $\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
@@
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma^{XX,i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}_{t}(s)$ back through the state impulse response $X_{t}(s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because the forward kernel environment is deterministic.
@@ A.3 The filtering kernel $F^{i}$
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u).$ (A.15)
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
@@ A.3.1 Derivation
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
@@
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
@@
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
@@
Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
The three terms admit a direct interpretation:
  • $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel. Nonzero only when $u\leq s$ (causal in the observation direction).
  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$. Nonzero only when $s\leq u$ (causal in the noise direction).
  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
Remark A.7 (Self-adjointness and uniqueness).
Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms exchange and the integral is invariant (since $P^{i}(r)$ is symmetric). Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map $(\mathcal{P}_{t}^{i}f)(s):=\Pi^{i}f(s)+\int_{0}^{t}F_{t}^{i}(r,s)^{\top}f(r)\,dr$ is self-adjoint, as required by its interpretation as an orthogonal projection. Uniqueness follows from Itô isometry: if $\tilde{F}$ also satisfies the density representation, then $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$ a.e.
@@
$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u),$
@@
Remark A.9 (Computational simplification).
The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration. In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
@@ A.3.2 Formal characterization
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
@@ Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
@@ B.2 Spike variation and the first-variation system
Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients—not a linearization but the full response. The deviator’s own innovation is unaffected ($\delta I_{s}^{i}=0$), since $\delta X_{s}\in\mathcal{F}_{t}^{i}$ implies $\delta\widehat{X}_{s}^{i}=\delta X_{s}$.
@@ B.6 Risk-sensitive extension: exponential utility FOC
The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure. Since $\mathcal{C}^{i}$ is quadratic in $W$ under the linear profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
B.7 Information cost: proof of Proposition 3.6
Full- and partial-information controls.
At the equilibrium profile $D^{*}$, the stochastic costate is
$M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u},$
where $(\bar{H}^{X,i},H^{X,i})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
$D^{i,\mathrm{full}}_{t}$ $=-(G^{DD,i}(t))^{-1}M^{i}_{t},$ (B.20)
$D^{i}_{t}$ $=-(G^{DD,i}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
Both are evaluated at the same adjoint kernels: opponents’ maps are fixed and linear with deterministic kernels, so the backward system is identical regardless of player $i$’s information. The control gap is
$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
where $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$.
Completion of the square.
Since $D^{i,\mathrm{full}}$ minimizes the strictly convex quadratic $D^{i}\mapsto J^{i}(D^{i};D^{*,-i})$ over $L^{2}(\mathcal{F})$ (Lemma 3.4), and $\delta D^{i}\in L^{2}(\mathcal{F})$, the first variation vanishes and
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{DD,i}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
To verify that only $G^{DD,i}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$. The zero-order term vanishes by definition. All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{X,i})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$. The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{DD,i}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal). The standard LQG completion-of-the-square identifies these: using $G^{DD,i}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
Kernel formula for $\Sigma^{M,i}_{t}$.
Substituting the control gap into (B.22):
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{DD,i}(t))^{-1}\left(\int_{0}^{t}H^{X,i}_{t}(v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{DD,i}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.7). It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
The conditional mean of $M^{i}_{t}$ given $\mathcal{F}^{i}_{t}$ is
$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}\widehat{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the resolved adjoint kernel is
$\widehat{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)\,\Pi^{i}+\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+\bigl(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s}\bigr)\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini. Then
$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the unresolved adjoint kernel $\widetilde{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)-\widehat{H}^{X,i}_{t}(s)=H^{X,i}_{t}(s)(I-\Pi^{i})-\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$ is the analogue of $\widetilde{X}^{i}$ for the backward system. By Itô isometry,
$\Sigma^{M,i}_{t}=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\widetilde{H}^{X,i}_{t}(s)^{\top}\,\mathrm{d}s.$
Separation from control.
The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$. The adjoint $H^{X,i}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself. Both $\widetilde{H}^{i}_{X}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof. ∎

### 23 Mar 2026 v3

@@ Appendix A Filtering
We identify the kernel equations for $F^{i}$ via formal calculations, then define $F^{i}$ as their unique solution (uniqueness follows from $L^{2}$-projection). The pipeline proceeds in three steps. Subsection A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations. Subsection A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain and establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$. Subsection A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
@@ A.2 Specialization to the physical observation model
$dY_{t}^{i}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},$
which corresponds to setting $C(t,s)=\Gamma^{i}(t)\,X_{t}(s)$ in the general framework of Proposition A.2.
@@
Substituting $C(t,s)=\Gamma^{i}(t)\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\Gamma^{i}(t)\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
@@
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\Gamma^{i}(t)\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\Gamma^{i}(t)(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$. Thus $\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
@@
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma^{XX,i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}_{t}(s)$ back through the state impulse response $X_{t}(s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because opponents’ strategy maps have deterministic coefficients.
@@ A.3 The filtering kernel $F^{i}$
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\Gamma^{i}(t)\,\widetilde{X}^{i}_{t}(u).$ (A.15)
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
@@ A.3.1 Derivation
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}\Bigg)\,du,$
@@
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\Gamma^{i}(u)\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
@@
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
@@
Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
The three terms have the same interpretation as in Proposition A.3.
@@
$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\Gamma^{i}(t)\,\widetilde{X}^{i}_{t}(u),$
@@ A.3.2 Formal characterization
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
@@ Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 4.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [19, Theorem 7.12]; Lemma 4.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 4.5.
@@ B.2 Spike variation and the first-variation system
Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients. The deviator’s own innovation is unaffected ($\delta I_{s}^{i}=0$), since $\delta X_{s}\in\mathcal{F}_{t}^{i}$ implies $\delta\widehat{X}_{s}^{i}=\delta X_{s}$.
@@ B.6 Risk-sensitive extension: exponential utility FOC
Appendix C Endogenous Signal Precision
This appendix identifies the obstruction to precision optimization in the multi-player setting.
C.1 Why precision variation is hard
In single-agent LQG, the separation principle makes precision optimization straightforward: the policy gain $S(t)$ does not depend on signal precision, so the value of information can be computed independently of the control problem. With multiple agents and endogenous signals, separation fails (Proposition 4.6): the policy kernel $D^{i}_{t}(u)$ depends on precision through the information wedge, so the marginal value of precision depends on the equilibrium.
A deeper difficulty concerns the observation model itself. The observation equation (3.4),
$dY^{i}_{t}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},$
treats each player’s information channel as exogenously given. This is adequate when the analyst asks: “holding the information structure fixed, what is the equilibrium?” It does not extend to a player choosing which sources to monitor, because rescaling the gain $\Gamma^{i}\to\Gamma^{i}+\delta\Gamma^{i}$ changes the filtration in a way that is not monotone.
Concretely, the filtrations generated by $\Gamma^{i}X\,dt+dW^{i}$ and $(\Gamma^{i}+\delta\Gamma^{i})X\,dt+dW^{i}$ are generically incomparable: neither contains the other. A player operating at the perturbed precision cannot reconstruct the noise-state path $\widehat{W}^{i}_{t}(\cdot)$ they would have maintained under the original gain—they have different information, not more information. This breaks the envelope argument that makes single-agent rational inattention tractable: the player at $P^{i}+\delta P^{i}$ cannot implement the control they would have chosen at $P^{i}$.
The incomparability disappears when precision is increased by adding an independent source (the Brownian-sheet formulation of Section 3.2). Monitoring an additional channel strictly enlarges the filtration, and the marginal value of the source is the value of optimally exploiting its independent increment. The full rational-inattention game—in which source portfolios are chosen endogenously—requires this formulation and is left to future work.
C.2 Two channels from a precision perturbation
Despite the incomparability, one can compute the first-order effect of a gain perturbation on equilibrium objects by linearizing the forward–backward system. The calculation reveals two channels through which $\delta P^{i}$ affects cost.
Fix an equilibrium profile and spike-perturb player $i$’s gain at time $t_{0}$. By the envelope theorem, $\delta D^{i}$ has no first-order cost effect (the FOC is satisfied). But the primitive-noise control $\mathcal{D}^{i}_{\tau}(r)=D^{i}_{\tau}(r)\Pi^{i}+\int_{0}^{\tau}D^{i}_{\tau}(u)\,F^{i}_{\tau}(u,r)\,du$ responds at first order through $\delta F^{i}$: the filtering kernel shifts, changing which part of the fixed policy kernel is resolved by the player’s own information.
The variation of the unresolved adjoint decomposes as
[…]
where $\mathcal{G}^{i}$ is the cost propagator and $\widetilde{\delta\mathcal{D}}^{i}$ is the unresolved component of the control variation induced by $\delta F^{i}$.
Channel (a) resolves more of the fixed adjoint—the standard value-of-information effect present in single-agent problems. Channel (b) arises because the same $\delta F^{i}$ shifts the projection boundary, pushing part of $\mathcal{D}^{i}_{\tau}$ outside the player’s current resolution. This is the equilibrium channel: it operates through the information wedge, and vanishes precisely when $\mathcal{V}^{i}_{t}=0$ (exogenous signals, Corollary 4.3). Multi-player rational inattention therefore couples the filtering and control problems that separate in single-agent settings: the marginal value of precision depends on the equilibrium through the wedge, and optimizing attention requires solving for the equilibrium first.